Two battery chargers (one for your laptop and the other for your phone) are plugged in at the same powerpoint.

a) If the phone charger consumes 30 W of electrical power when operating, how much current would flow through it when connected to a household supply? (Show your working and explain your answer.)

b) Explain how the laptop battery charger would be able to provide electrical safety to the operator in case of a short circuit. The laptop charger has three prongs while the mobile phone charger has two prongs.

c) Explain how the current and voltage will be distributed between the laptop and the phone chargers (Assume that the laptop battery charger has a higher resistance compared to the mobile phone charger).
Answer #1

a) P= V*I ( Here circuit is assumed to be resistive so of is 1)

I= P/V

I= 30/230

I= 130mA

b) In case of short circuit fault current will flow through earth pin of laptop charger to earth ensuring safety of personnel working and also safety of device.

c) Both circuit are in parallel connection so voltage will be same in both device. Laptop battery charger will take less current compare to mobile charger as it has higher resistance as current division rule in parallel circuit.